the second term of a geometric sequence is 10 and the fifth term is 7290. find the nth term
Answer by stanbon(75887)

the second term of a geometric sequence is 10 and the fifth term is 7290. find the nth term
-------------------
Form: a(n) = ab^(n-1)
----------------------
a(5) = ab^4 = 7290
a(2) = ab = 10
----
Divide a(5) by a(2) to solve for "b":
b^3 = 729
-----
b = 9
--------
Solve for "a":
ab = 10
a*9 = 10
a = 10/9 = 1.11111..
----
nth term = (10/9)(9)^n
a(n) = 10*9^(n-1)
========================
Cheers,
Stan H.